A computer can be classified as either cutting dash edge or ancient. Suppose that 85​% of computers are classified as ancient. ​(a) Two computers are chosen at random. What is the probability that both computers are ancient​? ​(b) Eight computers are chosen at random. What is the probability that all eight computers are ancient​? ​(c) What is the probability that at least one of eight randomly selected computers is cutting dash edge​? Would it be unusual that at least one of eight randomly selected computers is cutting dash edge​?
Answer #1

P = 0.85

It is a bimodal sldistribution.

P(X = x) = nCx * px * (1 - p)n - c

A) P(X = 2) = 2C2 * (0.85) ^2 * (0.15)^0 = 0.7225

B) P(X = 8) = 8C8 * (0.85)^8 * (0.15)^0 = 0.2725

C) P(X > 1) = 1 - P(X < 1)

= 1 - P(X = 0)

= 1 - (8C0 * (0.85)^0 * (0.15)^8)

= 1 - 0.0000 = 1

Since the probability value is greater than 0.95, so it would be unusual that at least one of eight randomly selected computers is cutting dash edge.